OpenStudy (anonymous):

if something is 148 dollars and i have a coupon for 30 dollars off 100 and i live in ohio how much is my total going to be with tax? tax is 7.75% @whpalmer4 @hedgehog @HELPME5846 @H265 @Preetha @OneOwOne @Xishem @Xetrevon @xblacem @UnkleRhaukus @uri @bobbylynn @beezz12 @beanz @faded911

OpenStudy (anonymous):

@damien @derek11111 @DDCamp @dan815 @radar @robtobey @RolyPoly @reemii @yulisasandoval98 @yoyogators @yentruok

OpenStudy (anonymous):

@Rebrox @RolyPoly @Olivia2424 @mlb004 @Mertsj @beezz12 @Xetrevon @aaronq @ajprincess @amistre64 @UnkleRhaukus @kristinaaaaa @khurramshahzad @kirbykirby @kyrsten17 @Poplar @pak_asfand @ParthKohli @i0iz0gangster @IrishBoy123 @iluvreading

OpenStudy (anonymous):

can someone help me please

OpenStudy (anonymous):

@Ell @yulisasandoval98 @OneOwOne @UnkleRhaukus @IrishBoy123 @i0iz0gangster @iluvreading @ivettef365 @hannah98xo @H265

OpenStudy (anonymous):

@Xishem please help me

OpenStudy (ell):

Sorry:( Math isn't my best subject. Hope you find the help you need! GBU!

OpenStudy (anonymous):

@GoldPhenoix @GodBlessYou @GrammarNazi @churrity

OpenStudy (anonymous):

@mathslover @mathisfun13 @mathkid55 @myko @khurramshahzad @kyrsten17 @kristinaaaaa

OpenStudy (anonymous):

please help me @GoldPhenoix

OpenStudy (goldphenoix):

I'm not sure if I can give you the correct answer.

OpenStudy (anonymous):

@thomaster @april115 @iamnotahuman

OpenStudy (anonymous):

@Luigi0210 @e.cociuba @MaxSteroids @surjithayer @sexygirl7400 @selenamalter

OpenStudy (anonymous):

can anyone help me please

OpenStudy (goldphenoix):

"coupon for 30 dollars off 100" is the part that is making me confused.

OpenStudy (anonymous):

the object is 146 dollar i have a coupon for 30 dollars of how much will be my total with tax the tax percentage is 7.75

OpenStudy (luigi0210):

Can't you make a proportion for that Gold?

OpenStudy (anonymous):

30 dollars off*************

OpenStudy (goldphenoix):

Oh. Now it make sense. (kinda) I'm thinking of doing 146 * 1.0775 - 30. Not sure if it's correct.

OpenStudy (anonymous):

from before i got 127.00

OpenStudy (anonymous):

is that right???????????

OpenStudy (anonymous):

130$

OpenStudy (anonymous):

sorry its 127.14 $

OpenStudy (anonymous):

so if it costs 148 dollars 30 off 100 would make it 118 dollars and then 7.75 % tax would make it 9.14+the cost (with 30%off) = $127.14

OpenStudy (anonymous):

cost of object=148$ coupon=30$ price=148-30=118$ tax=118*7.75=9.14$ total cost=118+9.14=127.14$

OpenStudy (anonymous):

Thats what i got!!

OpenStudy (anonymous):

lol me too.

OpenStudy (anonymous):

didn`nt get r question m not doing CA :P or CCA